Write an equation of the line containing the given point and perpendicular to the given line. Express your answer in the form yequalsmxplusb.
(3,9); 6x+y=5

Answer by Alan3354(69443)

Write an equation of the line containing the given point and perpendicular to the given line.
(3,9); 6x+y=5
-------------
Step 1, find the slope of the given line.
The slope of lines perpendicular is the negative inverse, call it m.
-------
Use y-y1 = m*(x-x1) where (x1,y1) is the point (3,9)
==============
This is basic. Learn it. Know it.